THE r_{0} STRUCTURAL PARAMETERS OF EQUATORIAL BROMOCYCLOBUTANE, CONFORMA-TIONAL STABILITY FROM TEMPERATURE DEPENDENT INFRARED SPECTRA OF XENON SOLU-TIONS, AND VIBRATIONAL ASSIGNMENTS

\underline{ARINDAM GANGULY}, JOSHUA J. KLAASSEN, TODOR K. GOUNEV, JAMES R. DURIG, DE-PARTMENT OF CHEMISTRY, UNIVERSITY OF MISSOURI-KANSAS CITY, KANSAS CITY, MO 64110,USA; GAMIL A. GUIRGIS, DEPARTMENT OF CHEMISTRY AND BIOCHEMISTRY, COLLEGE OF CHARLESTON, CHARLESTON, SC 29424, USA.

Variable temperature (-55 to −100^{∘}C) studies of the infrared spectra (4000-400 cm^{−1}) of bromocyclobutane, c-C_{4}H_{7}Br dissolved in liquid xenon have been carried out. The infrared spectrum (4000-100 cm^{−1}) of the gas has also been recorded. The enthalpy difference between the more stable equatorial conformer and the axial form, has been determined to be 372 ± 34 cm^{−1} . This experimental value of ΔH is much lower than the average MP2(full) ab initio predicted value of 521 ± 87 cm^{−1}. The percentage of the axial conformer present at ambient temperature is estimated to be 14 ± 1%. By utilizing previously reported microwave rotational constants for the equatorial conformer combined with ab initio MP2(full)/6-311+G(d,p) predicted structural values, adjusted r_{0} parameters have been obtained. The determined heavy atom structural parameters for this conformer are with distances(Å) C_{α}-Br = 1.942(3), C_{α}-C_{β} = 1.541(5), C_{γ}-C_{β} = 1.552(3)Å and angles in degrees ^{}C̸_{β}C_{α}Br = 118.4(5), ^{}C̸_{β}C_{α}C_{β} = 89.7(5), ^{}C̸_{γ}C_{β}C_{α} = 86.8(5), ^{}C̸_{β}C_{γ}C_{β} = 88.9(5) and τC_{γ}C_{β}C_{β}C_{α} = 29.8(5)^{∘}. The results will be discussed and compared to the corresponding properties of some similar molecules.